OpenStudy (anonymous):

do i need to find a common denominator if i'm subtracting two fractions??????

OpenStudy (gabebae):

yes

OpenStudy (alexandervonhumboldt2):

you need to make the denominators equal to subtartc fractions

OpenStudy (alexandervonhumboldt2):

you caanot subtract before you make denominators equal

OpenStudy (anonymous):

definitely XD

OpenStudy (anonymous):

ok so if i wanted to subtract 6/25 - 4/5 what common denominator would i use and y?

OpenStudy (alexandervonhumboldt2):

the common is 25

OpenStudy (alexandervonhumboldt2):

6/25 - 4/5=6/25 - 20/3=25=-14/25

OpenStudy (anonymous):

so i would use 25 as my common denominator?

OpenStudy (anonymous):

4/8 - 2/3 8*3= 24 4/8=12/24 2/3 = 16/24 12/24-16/24= 1/2 - 2/3 = -1/6

OpenStudy (markchernioglo):

yup

OpenStudy (anonymous):

mines was an example of what to do

OpenStudy (amistre64):

there is no mathical need to have common denominators;but the process does make working with fractions easier to do without technology.

OpenStudy (anonymous):

yep their right XD

OpenStudy (markchernioglo):

thats what im saying im just sitting here and agreeing lol. :D

OpenStudy (anonymous):

lol ok thanks guys :)

OpenStudy (markchernioglo):

good job for asking the question and understanding.

OpenStudy (anonymous):

thank you :)

OpenStudy (markchernioglo):

\(\Huge\cal\rlap{\color{red}{GREAT~JOB!}}{\color{blue}{\;GREAT~JOB!}}\)

OpenStudy (markchernioglo):

NP :)

OpenStudy (anonymous):

@markchernioglo XD

OpenStudy (markchernioglo):

lol